OpenStudy (anonymous):

spherical shell of inside radi. a and outside radi. b, b>a,has a spherically symmetric volume charge densi.ρ(r)=k/r² for ab. a)Deter. the elec.field strength at arbitrary points inside and outside the shell using Gauss' law. b)Integrating the results of part a inward from infinity, find the poten. at an arbitrary point within the shell (a

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Well.. to calculate the electric field at any point outside the shell is pretty easy.. so ll leave that up to you Inside the shell is zero.. cause no charge is enclosed.. but at any random point p at a distance r from the centre such that a<r<b..  apply Gauss law \[ \int\limits_{}^{}\bar E.\bar {dA}= \frac{Q_{enc}}{\epsilon_0}\] Due to symmetry.. the dot product vanishes, and you get E out.. and you can easily evaluate that trivial integral.. what now is Q_enc? well \[\sigma = \frac{dQ}{dv}\] or \[Q = \int\limits \sigma dv\] can u calculate further?!